OpenStudy (candy13106):

Evaluate the square root of -4 times the square root of -25 and the square root of 100. Does the rule square root of a times the square root of b = square root of ab hold when a and b are negative numbers.

OpenStudy (jdoe0001):

\(\bf \sqrt{-4}\cdot \sqrt{-25}\cdot \sqrt{100}?\)

OpenStudy (candy13106):



OpenStudy (jdoe0001):

so... do you undersand what " Does the rule square root of a times the square root of b = square root of ab hold when a and b are negative numbers. " means? that is\(\bf \sqrt{a}\cdot \sqrt{b}\iff\sqrt{a\cdot b}\) does that apply to negative radicands

OpenStudy (jdoe0001):

that is, do you understand what you're asked?

OpenStudy (candy13106):

yes

OpenStudy (jdoe0001):

hm

OpenStudy (candy13106):

ok

OpenStudy (jdoe0001):

\(\bf \sqrt{-4}\cdot \sqrt{-25}\implies \sqrt{-4\cdot -25}\implies \sqrt{+100}\) is that true?

OpenStudy (candy13106):

yea

OpenStudy (jdoe0001):

well, I have disappointing news :| is not \(\bf \sqrt{-4}\cdot \sqrt{-25}\implies\sqrt{-1\cdot 4}\cdot \sqrt{-1\cdot 25} \\ \quad \\ \sqrt{-1}\cdot \sqrt{4}\cdot \sqrt{-1}\cdot \sqrt{25}\implies 2i\cdot 5i\implies 10{\color{brown}{ i^2}}\implies 10({\color{brown}{ -1}})\implies -10\)

OpenStudy (candy13106):

ok

OpenStudy (jdoe0001):

that rule, does not apply to negative radicands, when the root is EVEN you first, take out the imaginary part, that is \(\bf \sqrt{-1}\) and then simplify

OpenStudy (candy13106):

ok thanks i understand it now

OpenStudy (jdoe0001):

yw